# Inference on testing the number of spikes in a high-dimensional generalized spiked Fisher matrix

### v2

@@ 3 Application to large-dimensional linear regression models
$\mathbf{z}_{i}=\mathbf{B}\mathbf{w}_{i}+\bm{\epsilon}_{i},\quad i=1,2,\ldots,n,$ (3.1)
@@
$\mathcal{H}_{0}:M=M_{0}\quad v.s.\quad\mathcal{H}_{1}:M\neq M_{0}.$ (3.2)
@@
${\Lambda}_{n}=\left|\mathbf{I}+\frac{r_{1}}{n-r}\mathbf{H}_{0}\mathbf{G}^{-1}\right|^{-1},$ (3.3)
@@
$(n-r)\mathbf{G}\sim W_{p}(\mathbf{V},n-r),\quad r_{1}\mathbf{H}_{0}\sim W_{p}(\mathbf{V},r_{1}),$ (3.4)
@@
To apply Theorem 2.1 to test the hypothesis (3.2), we first construct a spiked Fisher matrix with $M_{0}$ spikes. We start by decomposing $\mathbf{H}_{0}$ as $\mathbf{H}_{0}=(\hat{\mathbf{B}}_{1}\mathbf{A}_{11:2}\hat{\mathbf{B}}_{1}^{T}-\hat{\mathbf{B}}_{1}\mathbf{A}_{11:2}\mathbf{B}_{1}^{T}-\mathbf{B}_{1}\mathbf{A}_{11:2}\hat{\mathbf{B}}_{1}^{T}+\mathbf{B}_{1}\mathbf{A}_{11:2}\mathbf{B}_{1}^{T})/r_{1}$. Then define $\mathbf{H}=\hat{\mathbf{B}}_{1}\mathbf{A}_{11:2}\hat{\mathbf{B}}_{1}^{T}/r_{1}$, so that
@@
where $\mathbf{\Delta}_{0}=(\hat{\mathbf{B}}_{1}\mathbf{A}_{11:2}\mathbf{B}_{1}^{T}+\mathbf{B}_{1}\mathbf{A}_{11:2}\hat{\mathbf{B}}_{1}^{T}-\mathbf{B}_{1}\mathbf{A}_{11:2}\mathbf{B}_{1}^{T})/r_{1}$. Here, $\mathbf{\Delta}_{0}$ has $M_{0}$ dominated eigenvalues, indicating the spikes, and thus the matrix $\mathbf{H}$ represents a perturbation of rank $M_{0}$ on the matrix $\mathbf{H}_{0}$. We demonstrate the spiked structure and the eigenvalue properties of $\mathbf{H}$ in the supplementary material. Combining the formula (3.4), the matrix $\mathbf{H}\mathbf{G}^{-1}$ corresponds exactly to the spiked Fisher matrix with $M_{0}$ spikes defined in (2.2) under the null.
@@
$\Lambda_{n}^{*}=\left|\mathbf{I}+\frac{r_{1}}{n-r}\mathbf{H}\mathbf{G}^{-1}\right|^{-1}.$ (3.5)
Denote the eigenvalues of $\mathbf{H}\mathbf{G}^{-1}$ in (3.5) in descending order as below:
@@
For the testing problem (3.2) in the large-dimensional linear regression model in (3.1), $\Lambda_{n}^{*}$ is defined as in the formula (3.5), and it is assumed that $p/r_{1}\rightarrow c_{1}\in(0,\infty)$, $p/(n-r)\rightarrow c_{2}\in(0,1)$ as $\min(p,r_{1},n-r)\rightarrow\infty$. Then, under the null, we have
$T_{l}=\nu_{l}^{-\frac{1}{2}}\left[-\log{\Lambda}_{n}^{*}-\sum_{i=1}^{M}\log(1+\frac{c_{2}}{c_{1}}l_{i})-d_{l1}+d_{l2}-\mu_{l}\right]\Rightarrow\mathcal{N}(0,1),$ (3.6)
@@ 4 Application to change point detection
$\mathcal{H}_{0}:M_{j}=0\quad v.s.\quad\mathcal{H}_{1}:M_{j}\neq 0,j=1,2,\cdots.$ (4.1)
This is a degenerate hypothesis for the problem (2.6).
@@
To develop the CLT of the test statistic, we introduce some notations. Normalize the first group of data $\mathbf{X}_{j}^{(1)}$ as $\bm{\xi}_{j}^{(1)}$, and the second group of data $\mathbf{X}_{j}^{(2)}$ as $\bm{\xi}_{j}^{(2)}$. Further, define $\phi_{j}^{(1)}=E\left|\xi_{j,11}^{(1)}\right|^{4}$, $\phi_{j}^{(2)}=E\left|\xi_{j,11}^{(2)}\right|^{4}$, where $\xi_{j,11}^{(1)}$ and $\xi_{j,11}^{(2)}$ are the $(1,1)$-th entries of the matrices $\bm{\xi}_{j}^{(1)}$ and $\bm{\xi}_{j}^{(2)}$, respectively. Then, the asymptotic distribution of the test statistic $\mathrm{tr}[(\mathbf{S}_{j}^{(1)})^{-1}\mathbf{S}_{j}^{(2)}]$ is established as follows, which is a natural consequence of Theorem 2.1, and the calculation can refer to Example 2.1.
@@
For the testing problem (4.1) in each window $W_{j},j=1,2,\cdots$, assuming that $p/q_{j1}\rightarrow c_{j}^{(1)}\in(0,1)$, $p/q_{j2}\rightarrow c_{j}^{(2)}\in(0,+\infty)$ as $\min(p,q_{j1},q_{j2})\rightarrow\infty$. Then, under the null, we have
$T_{j}=\nu_{x}^{-\frac{1}{2}}\left\{\mathrm{tr}[(\mathbf{S}_{j}^{(1)})^{-1}\mathbf{S}_{j}^{(2)}]-d_{x}-\mu_{x}\right\}\Rightarrow\mathcal{N}(0,1),$ (4.2)
@@
5:     Calculate $T_{j}$ in the window $W_{j}$ according to (4.2);
@@ 5 Numerical and empirical studies
Tables 1 and 2 report the empirical sizes and powers of rejecting the null hypothesis (2.6) over 1,000 replications in Models 1 and 2 at a significance level $\alpha=0.05$ when $f(x)=\log(x)$. For space considerations, similar results for the function $f(x)=x$ are presented in the supplementary material. Under different models and population assumptions, the empirical sizes of our proposed test statistic are approximately equal to the significance level when the null hypothesis is true. The further the alternative hypothesis is from the null hypothesis, the higher the power is. From these results, it can be inferred that the true value $M_{0}$ aligns with the location of the first local minimum of empirical sizes. Furthermore, Figures 1 and 2 show the empirical density histograms of our proposed test statistic under the null for cases where $p=200$. These figures suggest that the proposed statistic is asymptotically normal when the null hypothesis is true, further validating the effectiveness of our method.
Table 1: Empirical size of rejecting the null hypothesis (2.6) in Model 1 when $f(x)=\log x$.
@@
Table 2: Empirical size of rejecting the null hypothesis (2.6) in Model 2 when $f(x)=\log x$.
@@ 5.2 Numerical studies for large-dimensional linear regression
For the hypothesis (3.2) in the large-dimensional linear regression model, we generate data from the following form:
@@ 6 Conclusion
$T=\sum\limits_{j=1}^{p}f\{l_{j}({\bf F})\}-\sum\limits_{j\in{\mathcal{J}}_{k},k=1}^{K}f\{l_{j}({\bf F})\}.$ (A.1)
We first focus on the calculation of the second term $\sum\limits_{j\in{\mathcal{J}}_{k},k=1}^{K}f\{l_{j}({\bf F})\}$ in (A.1). It is known that, for the sample spiked eigenvalues of $\mathbf{F}$, almost surely,
$\sum_{j\in\mathcal{J}_{k},k=1}^{K}f\{l_{j}({\bf F})\}\rightarrow\sum_{k=1}^{K}m_{k}f\left(\psi_{k}\right),$ (A.2)
where $\psi_{k}$ is defined in (2.5).
For the first term in (A.1), we have
@@
$\ +\frac{\beta_{y}c_{2}}{2\pi i}\oint f(z)\underline{m}^{\prime}(z)\frac{\underline{m}^{3}(z)m^{3}_{0}(z)h_{m_{2}}\{-\underline{m}^{-1}(z)\}}{1-c_{2}\int\frac{m^{2}_{0}(z)}{\big{\{}t+m_{0}(z)\big{\}}^{2}}{\rm d}H(t)}{\rm d}z,$ (A.3)
@@
$\quad\cdot\frac{\partial^{2}\left[\underline{m}\left(z_{1}\right)\!m_{0}\left(z_{1}\right)\!\underline{m}\left(z_{2}\right)\!m_{0}\left(z_{2}\right)\!h_{v_{2}}\left(-\underline{m}^{-1}\left(z_{1}\right),\!-\underline{m}^{-1}\left(z_{2}\right)\right)\right]}{\partial z_{1}\partial z_{2}}{\rm d}z_{1}{\rm d}z_{2}.$ (A.4)
@@
The contours all contain the support of the LSD of ${\bf F}$ and are non-overlapping in (A.3) and (A.4).
@@
The authors are grateful to the Editor, the Associate Editors, and the referees for their review of the paper. The second author is supported by Key technologies for coordination and interoperation of power distribution service resource, Grant No. 2021YFB2401300 and NSFC Grant No.11971371.
@@
Supplementary materials contain an algorithm to approximate the center parameter term $d_{1}(f,H_{n})$ in Theorem 2.1, calculation details of Examples 2.1 and 2.2, an illustration about properties of the matrix $\mathbf{H}$ in Section 3, the additional simulation results for Section 5.1, and details of the selection method of the significance level $\alpha$ in Section 5.3.

### v3

@@ 3 Application to large-dimensional linear regression models
$\mathbf{z}_{i}=\mathbf{B}\mathbf{w}_{i}+\bm{\epsilon}_{i},\quad i=1,2,\ldots,n,$ (13)
@@
$\mathcal{H}_{0}:M=M_{0}\quad v.s.\quad\mathcal{H}_{1}:M\neq M_{0}.$ (14)
@@
${\Lambda}_{n}=\left|\mathbf{I}+\frac{r_{1}}{n-r}\mathbf{H}_{0}\mathbf{G}^{-1}\right|^{-1},$ (15)
@@
$(n-r)\mathbf{G}\sim W_{p}(\mathbf{V},n-r),\quad r_{1}\mathbf{H}_{0}\sim W_{p}(\mathbf{V},r_{1}),$ (16)
@@
To apply Theorem 2.1 to test the hypothesis (14), we first construct a spiked Fisher matrix with $M_{0}$ spikes. We start by decomposing $\mathbf{H}_{0}$ as $\mathbf{H}_{0}=(\hat{\mathbf{B}}_{1}\mathbf{A}_{11:2}\hat{\mathbf{B}}_{1}^{T}-\hat{\mathbf{B}}_{1}\mathbf{A}_{11:2}\mathbf{B}_{1}^{T}-\mathbf{B}_{1}\mathbf{A}_{11:2}\hat{\mathbf{B}}_{1}^{T}+\mathbf{B}_{1}\mathbf{A}_{11:2}\mathbf{B}_{1}^{T})/r_{1}$. Then define $\mathbf{H}=\hat{\mathbf{B}}_{1}\mathbf{A}_{11:2}\hat{\mathbf{B}}_{1}^{T}/r_{1}$, so that
@@
where $\mathbf{\Delta}_{0}=(\hat{\mathbf{B}}_{1}\mathbf{A}_{11:2}\mathbf{B}_{1}^{T}+\mathbf{B}_{1}\mathbf{A}_{11:2}\hat{\mathbf{B}}_{1}^{T}-\mathbf{B}_{1}\mathbf{A}_{11:2}\mathbf{B}_{1}^{T})/r_{1}$. Here, $\mathbf{\Delta}_{0}$ has $M_{0}$ dominated eigenvalues, indicating the spikes, and thus the matrix $\mathbf{H}$ represents a perturbation of rank $M_{0}$ on the matrix $\mathbf{H}_{0}$. We demonstrate the spiked structure and the eigenvalue properties of $\mathbf{H}$ in the supplementary material. Combining the formula (16), the matrix $\mathbf{H}\mathbf{G}^{-1}$ corresponds exactly to the spiked Fisher matrix with $M_{0}$ spikes defined in (4) under the null.
@@
$\Lambda_{n}^{*}=\left|\mathbf{I}+\frac{r_{1}}{n-r}\mathbf{H}\mathbf{G}^{-1}\right|^{-1}.$ (17)
Denote the eigenvalues of $\mathbf{H}\mathbf{G}^{-1}$ in (17) in descending order as below:
@@
For the testing problem (14) in the large-dimensional linear regression model in (13), $\Lambda_{n}^{*}$ is defined as in the formula (17), and it is assumed that $p/r_{1}\rightarrow c_{1}\in(0,\infty)$, $p/(n-r)\rightarrow c_{2}\in(0,1)$ as $\min(p,r_{1},n-r)\rightarrow\infty$. Then, under the null, we have
$T_{l}=\nu_{l}^{-\frac{1}{2}}\left[-\log{\Lambda}_{n}^{*}-\sum_{i=1}^{M}\log(1+\frac{c_{2}}{c_{1}}l_{i})-d_{l1}+d_{l2}-\mu_{l}\right]\Rightarrow\mathcal{N}(0,1),$ (18)
@@ 4 Application to change point detection
$\mathcal{H}_{0}:M_{j}=0\quad v.s.\quad\mathcal{H}_{1}:M_{j}\neq 0,j=1,2,\cdots.$ (19)
This is a degenerate hypothesis for the problem (8).
@@
To develop the CLT of the test statistic, we introduce some notations. Normalize the first group of data $\mathbf{X}_{j}^{(1)}$ as $\bm{\xi}_{j}^{(1)}$, and the second group of data $\mathbf{X}_{j}^{(2)}$ as $\bm{\xi}_{j}^{(2)}$. Further, define $\phi_{j}^{(1)}=E\left|\xi_{j,11}^{(1)}\right|^{4}$, $\phi_{j}^{(2)}=E\left|\xi_{j,11}^{(2)}\right|^{4}$, where $\xi_{j,11}^{(1)}$ and $\xi_{j,11}^{(2)}$ are the $(1,1)$-th entries of the matrices $\bm{\xi}_{j}^{(1)}$ and $\bm{\xi}_{j}^{(2)}$, respectively. Then, the asymptotic distribution of the test statistic $\mathrm{tr}[(\mathbf{S}_{j}^{(1)})^{-1}\mathbf{S}_{j}^{(2)}]$ is established as follows, which is a natural consequence of Theorem 2.1, and the calculation can refer to Example 1.
@@
For the testing problem (19) in each window $W_{j},j=1,2,\cdots$, assuming that $p/q_{j1}\rightarrow c_{j}^{(1)}\in(0,1)$, $p/q_{j2}\rightarrow c_{j}^{(2)}\in(0,+\infty)$ as $\min(p,q_{j1},q_{j2})\rightarrow\infty$. Then, under the null, we have
$T_{j}=\nu_{x}^{-\frac{1}{2}}\left\{\mathrm{tr}[(\mathbf{S}_{j}^{(1)})^{-1}\mathbf{S}_{j}^{(2)}]-d_{x}-\mu_{x}\right\}\Rightarrow\mathcal{N}(0,1),$ (20)
@@
5:     Calculate $T_{j}$ in the window $W_{j}$ according to (20);
@@ 5 Numerical and empirical studies
Tables 1 and 2 report the empirical sizes and powers of rejecting the null hypothesis (8) over 1,000 replications in Models 1 and 2 at a significance level $\alpha=0.05$ when $f(x)=\log(x)$. For space considerations, similar results for the function $f(x)=x$ are presented in the supplementary material. Under different models and population assumptions, the empirical sizes of our proposed test statistic are approximately equal to the significance level when the null hypothesis is true. The further the alternative hypothesis is from the null hypothesis, the higher the power is. From these results, it can be inferred that the true value $M_{0}$ aligns with the location of the first local minimum of empirical sizes. Furthermore, Figures 1 and 2 show the empirical density histograms of our proposed test statistic under the null for cases where $p=200$. These figures suggest that the proposed statistic is asymptotically normal when the null hypothesis is true, further validating the effectiveness of our method.
Table 1: Empirical size of rejecting the null hypothesis (8) in Model 1 when $f(x)=\log x$.
@@
Table 2: Empirical size of rejecting the null hypothesis (8) in Model 2 when $f(x)=\log x$.
@@ 5.2 Numerical studies for large-dimensional linear regression
For the hypothesis (14) in the large-dimensional linear regression model, we generate data from the following form:
@@ 6 Conclusion
$T=\sum\limits_{j=1}^{p}f\{l_{j}({\bf F})\}-\sum\limits_{j\in{\mathcal{J}}_{k},k=1}^{K}f\{l_{j}({\bf F})\}.$ (21)
We first focus on the calculation of the second term $\sum\limits_{j\in{\mathcal{J}}_{k},k=1}^{K}f\{l_{j}({\bf F})\}$ in (21). It is known that, for the sample spiked eigenvalues of $\mathbf{F}$, almost surely,
$\sum_{j\in\mathcal{J}_{k},k=1}^{K}f\{l_{j}({\bf F})\}\rightarrow\sum_{k=1}^{K}m_{k}f\left(\psi_{k}\right),$ (22)
where $\psi_{k}$ is defined in (7).
For the first term in (21), we have
@@
$\ +\frac{\beta_{y}c_{2}}{2\pi i}\oint f(z)\underline{m}^{\prime}(z)\frac{\underline{m}^{3}(z)m^{3}_{0}(z)h_{m_{2}}\{-\underline{m}^{-1}(z)\}}{1-c_{2}\int\frac{m^{2}_{0}(z)}{\big{\{}t+m_{0}(z)\big{\}}^{2}}{\rm d}H(t)}{\rm d}z,$ (23)
@@
$\quad\cdot\frac{\partial^{2}\left[\underline{m}\left(z_{1}\right)\!m_{0}\left(z_{1}\right)\!\underline{m}\left(z_{2}\right)\!m_{0}\left(z_{2}\right)\!h_{v_{2}}\left(-\underline{m}^{-1}\left(z_{1}\right),\!-\underline{m}^{-1}\left(z_{2}\right)\right)\right]}{\partial z_{1}\partial z_{2}}{\rm d}z_{1}{\rm d}z_{2}.$ (24)
@@
The contours all contain the support of the LSD of ${\bf F}$ and are non-overlapping in (23) and (24).
@@
{acks}
[Acknowledgments] The authors would like to thank the anonymous referees, the Associate Editors, and the Editor for their constructive comments of this paper.
@@
The second author was supported by Key technologies for coordination and interoperation of power distribution service resource Grant No. 2021YFB2401300, and National Natural Science Foundation of China Grant No. 12326606.
@@
Supplementary materials contain an algorithm to approximate the center parameter term $d_{1}(f,H_{n})$ in Theorem 2.1, calculation details of Examples 1 and 2, an illustration about properties of the matrix $\mathbf{H}$ in Section 3, the additional simulation results for Section 5.1, and details of the selection method of the significance level $\alpha$ in Section 5.3.